Solve the problem making a chart.
A Roman statue is three times as old as Florentine statue. One hundred years from now the Roman statue will be twice as old as Florentine statue. How old is the Roman statue?

Answer by ikleyn(52778)

.
First phrase says

    R = 3F                    (1)     (regarding ages)


Second phrase says


    R + 100 = 2*(F + 100).    (2)     (100 years after . . . )


Substitute (1) into (2) to get


    3F + 100 = 2F + 200


and solve


    3F - 2F = 100

    F = 100.


Answer.  F = 100;  R = 300.

-----------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - A fresh formulation of a traditional age problem
    - Really intricate age word problem
    - Selected age word problems from the archive
    - Age problems for mental solution
in this site.

Read them and become an expert in solving age problems.

Also, you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.